OpenStudy (anonymous):

A bag contains 5 black marbles and 4 white marbles. A spinner has 3 equal sections labeled 1 through 3. What is the probability of drawing a black marble from the bag and spinning a 1 on the spinner? A.5/9 B.1/2 C.2/9 D.5/27

OpenStudy (kirbykirby):

\[ P(\text{black})=\frac{5}{9}\]\[P(\text{1 on spinner})=\frac{1}{3} \] \[P(\text{black AND 1 on spinner})=P(\text{black})\cdot P(\text{1 on spinner}) \]